# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### 05 Mar 2026 v3

@@ 1.2 Our results: Average-case hardness of shallow-depth Boson Sampling
To sum up, we show the average-case hardness over outcomes and circuit instances for shallow circuit architectures composed of geometrically non-local gates and employing the local random circuit ensemble. We informally present here our average-case hardness result of Boson Sampling in the logarithmic depth regime, for photon number NN and mode number $M\propto N^{\gamma}$ with $\gamma\geq 2$.
@@
Consider a Boson Sampling problem with NN photons over $M=\Omega(N^{\gamma})$ mode, where $\gamma\geq 2$. There exists an $O(\log N)$-depth linear optical circuit architecture $\mathcal{A}$, consisting of $O(N^{\gamma}\log N)$ number of geometrically non-local gates, such that estimating the output probability to within additive error $2^{-O(N^{\gamma+1}(\log N)^{2})}$, with high probability over the choice of outcomes and circuits in $\mathcal{A}$, is $\rm{P}$-hard under $\rm{BPP}^{\rm{NP}}$ reduction.
@@ 2.1 Notations
Let us define the total mode number as $M$, where we set $M$ as a power of 2 for simplicity. We set the output photon number NN polynomially related to $M$ as $M=c_{0}N^{\gamma}$ for a constant $c_{0}$ and $\gamma\geq 2$, to ensure that collision-free outcomes dominate the probability weight over all possible NN-photon outcomes over $M$ modes (i.e., by bosonic birthday paradox; see Refs. [aaronson2011computational, arkhipov2012bosonic] and Appendix B). We use the notation $\bm{s}$ as an $M$-dimensional output configuration vector representing NN-photon outcome over $M$ modes, such that each element $s_{i}$ ($i\in[M]$) of $\bm{s}$ denotes the occupied number of photons in the $i$th mode. Namely, $\bm{s}$ can be expressed as $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in[N]$ with the constraint $\sum_{i=1}^{M}s_{i}=N$. Also, let $\bm{t}$ be an $M$-dimensional input configuration vector corresponding to an NN-photon input state over $M$ modes, such that $\bm{t}=(t_{1},\dots,t_{M})$ where now each $t_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}t_{i}=N$.
@@
where $C_{\bm{s},\bm{t}}$ is an NN by NN matrix obtained by taking $s_{i}$ copies of the $i$th row and $t_{j}$ copies of the $j$th column of the matrix $C$. Here, we omit the $\bm{t}$ dependence in the output probability $p_{\bm{s}}(C)$ for simplicity because we will fix the input configuration throughout the paper.
@@
Here, the circuit depth of the $q$-Kaleidoscope architecture is $D=2q\log M$, which is indeed a logarithmic depth $D=O(\log N)$ for $q=O(1)$. Also, the number of gates for the $q$-Kaleidoscope architecture is $m=qM\log M=O(N^{\gamma}\log N)$ for $q=O(1)$. Note that, because all the gates are geometrically non-local except for the gates in the first and the last depth for each $\mathcal{B}\mathcal{B}^{*}$, the number of non-local gates also scales $O(N^{\gamma}\log N)$ for $q=O(1)$. Throughout this work, we will focus on the $q$-Kaleidoscope circuit architecture with $q=O(1)$ to demonstrate the hardness results for shallow-depth circuits. The motivation for employing this linear optical circuit architecture is that it enjoys a useful property that is crucial for our analysis, which is that for $M$ a power of 22, any $M$-mode permutation circuit can be implemented within $\mathcal{B}\mathcal{B}^{*}$:
@@ 2.2 Definitions for random circuit and outcome ensemble
Note that this random permutation circuit ensemble can be implemented in the shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$ by Lemma 1. Before proceeding, we briefly remark on the motivation for including the permutation circuit ensemble in our random circuit construction. Specifically, for any $M$-mode circuit architecture $\mathcal{A}$, including $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ considered in this work, a circuit drawn from the product circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ guarantees the bosonic birthday paradox [aaronson2011computational, arkhipov2012bosonic]. The bosonic birthday paradox ensures that the probability of obtaining only collision outcomes is sufficiently large given that the photon number NN is sufficiently smaller than the mode number $M$, as stated in the following lemma.
@@
Let $M\geq 2N^{2}$, and let $B_{M,N}$ be a set of collision outcomes for $M$ modes and NN photons. Then, for a randomly chosen unitary circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$, the probability to obtain collision outcomes is upper bounded by
@@
We define $\mathcal{G}_{M,N}$ as the uniform distribution over $\binom{M}{N}$ number of collision-free outcomes of Boson Sampling with $M$ modes and NN photons. Each outcome $\bm{s}\sim\mathcal{G}_{M,N}$ is an $M$-dimensional output configuration vector representing the collision-free outcome, such that $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}s_{i}=N$.
@@ 6 Average-case hardness for shallow-depth Gaussian Boson Sampling
We also define an $M$-dimensional vector $\bm{s}$ as a serial concatenation of two $\bm{s}_{0}$ vectors, so that $\bm{s}$ represents an NN-photon outcome for $M$ modes. Then the output probabilities $p_{\bm{s}_{0}}(C_{0})$ and $q_{\bm{s}}(C)$ are related as
@@
Using the results of Theorem 6 and the previous proof of the average-case hardness of Boson Sampling in Theorem 4, it is straightforward to find the average-case hardness of Gaussian Boson Sampling, for a randomly chosen NN-photon outcomes $\bm{s}\sim\mathcal{G}_{M,N}$ and a randomly chosen circuit $U\sim\mathcal{H}_{\mathcal{A}}$ in shallow-depth architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$.
@@
where $\mu$ is along all possible perfect matching permutations over NN modes. From the proof of Lemma 10, using reduction to the common denominator for all of the $m=qM\log M$ gates, $\left[V(\theta)\right]_{j,k}$ can be represented as $(2m,2m)$ rational function in $\theta$ with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)$. Using this fact, one can easily check that $\prod_{j=1}^{N/2}\left[(V(\theta)V(\theta)^{T})_{\bm{s}}\right]_{\mu(2j-1),\mu(2j)}$ can be represented as $(2mN,2mN)$ rational function in $\theta$, with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)]^{N}$ which does not change with $\mu$. Therefore, the output probability can be represented as $q_{\bm{s}}(V(\theta))=\frac{P(\theta)}{Q(\theta)}$, with each $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$ and $P(\theta)$ a degree $4mN$ polynomial function in $\theta$.
@@ 7 Hardness analyses in lossy environments
To simplify, we assume that we know $\it{a\,priori}$ each error rate $\rho_{i}$ for all $i\in[l]$, and the noise model $\mathcal{N}$ is fixed so that it does not change with random circuit instances. Then we can obtain the hardness of approximating output probabilities of lossy shallow circuits, from our previous hardness proposals. To do so, let $p_{\bm{s}}(C,\mathcal{N})$ be the output probability of an NN-photon outcome $\bm{s}$ from an $M$-mode linear optical circuit $C$ which undergoes loss model $\mathcal{N}$ we set. By post-selecting ‘no loss event’, which can be accomplished by counting the output photon number, the ideal output probability $p_{\bm{s}}(C)$ can be inferred from $p_{\bm{s}}(C,\mathcal{N})$ by
@@
We remark that for our noise model, the imprecision blowup grows exponentially with the gate number $m$. Therefore, shallow-depth circuits can be more advantageous in this perspective, since they are likely to have less gate number and thus have small imprecision blowup. For example, the current hardness results are based on $M$ by NN submatrices of $M$-dimensional Haar random unitaries, and the implementation of such matrices requires gate number $m=\Omega(N^{\gamma+1})$. This arouses the imprecision blowup at least $2^{O(N^{\gamma+1})}$, which restricts the allowed additive error for the average-case hardness at most $2^{-O(N^{\gamma+1})}$.
@@ 8 Concluding remarks
Another possible approach for reducing the imprecision gap is to perturb a random circuit matrix in a different way from the Cayley transform (i.e., Definition 8), i.e., as depicted in [bouland2023complexity]. Specifically, instead of perturbing each random gate, one can perturb a submatrix $X$ of our random circuit matrix $U$ with a worst-case matrix AA as $X(\theta)=(1-\theta)X+\theta A$ for $\theta\in[0,1]$. Here, a degree of polynomial $|\text{Per}(X(\theta))|^{2}$ is $2N$, which is lower than ours derived by the Cayley transform. Therefore, if one can prove that $X(\theta)$ is distributed similarly to $X$ for small $\theta$, we expect that we can also obtain a better imprecision level by using the same interpolation method. However, the above approach requires one to figure out a global circuit distribution generated by the convolution of local circuit distributions. Although we believe that this problem can be resolved using techniques from random matrix theories, we have not yet developed a complete analysis. Hence, we leave it as an open question.
@@
3. One of the existing gaps between theoretical Boson Sampling hardness arguments and Boson Sampling experiments to date is the ratio between mode number $M$ and photon number NN. Generally, most of the current theoretical Boson Sampling hardness arguments (including our current work) consider only collision-free outcomes and rule out collision outcomes due to the convenience of mathematical analysis [aaronson2011computational, deshpande2022quantum, grier2022complexity, bouland2022noise, bouland2024average]. This calls for the “dilute" regime $M=\Omega(N^{2})$ to ensure the dominance of collision-free outcomes over the probability weight. In contrast, all the experiments to date [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian] operate in the “saturated regime” where the number of modes is comparable to the number of photons. Therefore, to reduce this gap, extending our hardness result of shallow-depth Boson Sampling to the saturated regime $M=O(N^{\gamma})$ for $1\leq\gamma<2$ is crucial, similarly to [bouland2023complexity] for the general Boson Sampling case. To do so, one also needs to consider collision outcomes in the analysis because collision outcomes now dominate the probability weight in this regime. Here, given that our worst-case hardness result in Theorem 3 holds for a linear number of modes $M=\Omega(N)$, this opens the possibility of establishing average-case hardness in the saturated regime by extending the worst-to-average-case reduction from Theorem 3 to include collision outcomes. However, since the collision outcomes do not interchange with each other generally by permutations (in contrast to collision-free outcomes), our proof technique in Sec. 4 using random permutation does not work for general collision outcomes. Therefore, more advanced proof techniques are required to address this challenge. In Appendix H, we present an approach that could offer a useful insight toward extending our average-case hardness result to include collision outcomes. Also, another possible way is to analytically prove that the local random circuit for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ is close enough to those of global Haar random circuits (numerically shown in [go2024exploring] and Appendix I). If this can be done, we can employ the previous argument [bouland2023complexity], which used global Haar random circuits, for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ in the saturated regime.
@@ Appendix C Proof of Theorem 3 (worst-case hardness)
We prove Lemma 7 by showing that the constant-depth linear optical circuit $C_{0}$ in Lemma 5 can be embedded into a 33-dimensional grid-structured circuit, which can be implemented in the $\mathcal{B}\mathcal{B}^{*}$ architecture under a suitable permutation of mode indices. For interested readers, we provide the full proof of Lemma 7 in Appendix D.
@@ D.2 Encoding a grid linear optical circuit to the butterfly circuit architecture $\mathcal{B}$
Suppose that a mode number is given by $M=2^{n}$ for $n\in\mathbb{Z}^{+}$. For an $M$-mode single parallel 11-dimensional grid linear optical circuit $C$, there exist an $M$-mode circuit $C^{\prime}$ in $\mathcal{B}$ and an $M$-mode permutation matrix $\bm{P}$ that satisfy $C=\bm{P}C^{\prime}\bm{P}^{T}$

### 06 Mar 2026 v4

@@ 1.2 Our results: Average-case hardness of shallow-depth Boson Sampling
To sum up, we show the average-case hardness over outcomes and circuit instances for shallow circuit architectures composed of geometrically non-local gates and employing the local random circuit ensemble. We informally present here our average-case hardness result of Boson Sampling in the logarithmic depth regime, for photon number $N$ and mode number $M\propto N^{\gamma}$ with $\gamma\geq 2$.
@@
Consider a Boson Sampling problem with $N$ photons over $M=\Omega(N^{\gamma})$ mode, where $\gamma\geq 2$. There exists an $O(\log N)$-depth linear optical circuit architecture $\mathcal{A}$, consisting of $O(N^{\gamma}\log N)$ number of geometrically non-local gates, such that estimating the output probability to within additive error $2^{-O(N^{\gamma+1}(\log N)^{2})}$, with high probability over the choice of outcomes and circuits in $\mathcal{A}$, is $\rm{P}$-hard under $\rm{BPP}^{\rm{NP}}$ reduction.
@@ 2.1 Notations
Let us define the total mode number as $M$, where we set $M$ as a power of 2 for simplicity. We set the output photon number $N$ polynomially related to $M$ as $M=c_{0}N^{\gamma}$ for a constant $c_{0}$ and $\gamma\geq 2$, to ensure that collision-free outcomes dominate the probability weight over all possible $N$-photon outcomes over $M$ modes (i.e., by bosonic birthday paradox; see Refs. [aaronson2011computational, arkhipov2012bosonic] and Appendix B). We use the notation $\bm{s}$ as an $M$-dimensional output configuration vector representing $N$-photon outcome over $M$ modes, such that each element $s_{i}$ ($i\in[M]$) of $\bm{s}$ denotes the occupied number of photons in the $i$th mode. Namely, $\bm{s}$ can be expressed as $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in[N]$ with the constraint $\sum_{i=1}^{M}s_{i}=N$. Also, let $\bm{t}$ be an $M$-dimensional input configuration vector corresponding to an $N$-photon input state over $M$ modes, such that $\bm{t}=(t_{1},\dots,t_{M})$ where now each $t_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}t_{i}=N$.
@@
where $C_{\bm{s},\bm{t}}$ is an $N$ by $N$ matrix obtained by taking $s_{i}$ copies of the $i$th row and $t_{j}$ copies of the $j$th column of the matrix $C$. Here, we omit the $\bm{t}$ dependence in the output probability $p_{\bm{s}}(C)$ for simplicity because we will fix the input configuration throughout the paper.
@@
Here, the circuit depth of the $q$-Kaleidoscope architecture is $D=2q\log M$, which is indeed a logarithmic depth $D=O(\log N)$ for $q=O(1)$. Also, the number of gates for the $q$-Kaleidoscope architecture is $m=qM\log M=O(N^{\gamma}\log N)$ for $q=O(1)$. Note that, because all the gates are geometrically non-local except for the gates in the first and the last depth for each $\mathcal{B}\mathcal{B}^{*}$, the number of non-local gates also scales $O(N^{\gamma}\log N)$ for $q=O(1)$. Throughout this work, we will focus on the $q$-Kaleidoscope circuit architecture with $q=O(1)$ to demonstrate the hardness results for shallow-depth circuits. The motivation for employing this linear optical circuit architecture is that it enjoys a useful property that is crucial for our analysis, which is that for $M$ a power of $2$, any $M$-mode permutation circuit can be implemented within $\mathcal{B}\mathcal{B}^{*}$:
@@ 2.2 Definitions for random circuit and outcome ensemble
Note that this random permutation circuit ensemble can be implemented in the shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$ by Lemma 1. Before proceeding, we briefly remark on the motivation for including the permutation circuit ensemble in our random circuit construction. Specifically, for any $M$-mode circuit architecture $\mathcal{A}$, including $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ considered in this work, a circuit drawn from the product circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ guarantees the bosonic birthday paradox [aaronson2011computational, arkhipov2012bosonic]. The bosonic birthday paradox ensures that the probability of obtaining only collision outcomes is sufficiently large given that the photon number $N$ is sufficiently smaller than the mode number $M$, as stated in the following lemma.
@@
Let $M\geq 2N^{2}$, and let $B_{M,N}$ be a set of collision outcomes for $M$ modes and $N$ photons. Then, for a randomly chosen unitary circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$, the probability to obtain collision outcomes is upper bounded by
@@
We define $\mathcal{G}_{M,N}$ as the uniform distribution over $\binom{M}{N}$ number of collision-free outcomes of Boson Sampling with $M$ modes and $N$ photons. Each outcome $\bm{s}\sim\mathcal{G}_{M,N}$ is an $M$-dimensional output configuration vector representing the collision-free outcome, such that $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}s_{i}=N$.
@@ 6 Average-case hardness for shallow-depth Gaussian Boson Sampling
We also define an $M$-dimensional vector $\bm{s}$ as a serial concatenation of two $\bm{s}_{0}$ vectors, so that $\bm{s}$ represents an $N$-photon outcome for $M$ modes. Then the output probabilities $p_{\bm{s}_{0}}(C_{0})$ and $q_{\bm{s}}(C)$ are related as
@@
Using the results of Theorem 6 and the previous proof of the average-case hardness of Boson Sampling in Theorem 4, it is straightforward to find the average-case hardness of Gaussian Boson Sampling, for a randomly chosen $N$-photon outcomes $\bm{s}\sim\mathcal{G}_{M,N}$ and a randomly chosen circuit $U\sim\mathcal{H}_{\mathcal{A}}$ in shallow-depth architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$.
@@
where $\mu$ is along all possible perfect matching permutations over $N$ modes. From the proof of Lemma 10, using reduction to the common denominator for all of the $m=qM\log M$ gates, $\left[V(\theta)\right]_{j,k}$ can be represented as $(2m,2m)$ rational function in $\theta$ with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)$. Using this fact, one can easily check that $\prod_{j=1}^{N/2}\left[(V(\theta)V(\theta)^{T})_{\bm{s}}\right]_{\mu(2j-1),\mu(2j)}$ can be represented as $(2mN,2mN)$ rational function in $\theta$, with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)]^{N}$ which does not change with $\mu$. Therefore, the output probability can be represented as $q_{\bm{s}}(V(\theta))=\frac{P(\theta)}{Q(\theta)}$, with each $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$ and $P(\theta)$ a degree $4mN$ polynomial function in $\theta$.
@@ 7 Hardness analyses in lossy environments
To simplify, we assume that we know $\it{a\,priori}$ each error rate $\rho_{i}$ for all $i\in[l]$, and the noise model $\mathcal{N}$ is fixed so that it does not change with random circuit instances. Then we can obtain the hardness of approximating output probabilities of lossy shallow circuits, from our previous hardness proposals. To do so, let $p_{\bm{s}}(C,\mathcal{N})$ be the output probability of an $N$-photon outcome $\bm{s}$ from an $M$-mode linear optical circuit $C$ which undergoes loss model $\mathcal{N}$ we set. By post-selecting ‘no loss event’, which can be accomplished by counting the output photon number, the ideal output probability $p_{\bm{s}}(C)$ can be inferred from $p_{\bm{s}}(C,\mathcal{N})$ by
@@
We remark that for our noise model, the imprecision blowup grows exponentially with the gate number $m$. Therefore, shallow-depth circuits can be more advantageous in this perspective, since they are likely to have less gate number and thus have small imprecision blowup. For example, the current hardness results are based on $M$ by $N$ submatrices of $M$-dimensional Haar random unitaries, and the implementation of such matrices requires gate number $m=\Omega(N^{\gamma+1})$. This arouses the imprecision blowup at least $2^{O(N^{\gamma+1})}$, which restricts the allowed additive error for the average-case hardness at most $2^{-O(N^{\gamma+1})}$.
@@ 8 Concluding remarks
Another possible approach for reducing the imprecision gap is to perturb a random circuit matrix in a different way from the Cayley transform (i.e., Definition 8), i.e., as depicted in [bouland2023complexity]. Specifically, instead of perturbing each random gate, one can perturb a submatrix $X$ of our random circuit matrix $U$ with a worst-case matrix $A$ as $X(\theta)=(1-\theta)X+\theta A$ for $\theta\in[0,1]$. Here, a degree of polynomial $|\text{Per}(X(\theta))|^{2}$ is $2N$, which is lower than ours derived by the Cayley transform. Therefore, if one can prove that $X(\theta)$ is distributed similarly to $X$ for small $\theta$, we expect that we can also obtain a better imprecision level by using the same interpolation method. However, the above approach requires one to figure out a global circuit distribution generated by the convolution of local circuit distributions. Although we believe that this problem can be resolved using techniques from random matrix theories, we have not yet developed a complete analysis. Hence, we leave it as an open question.
@@
3. One of the existing gaps between theoretical Boson Sampling hardness arguments and Boson Sampling experiments to date is the ratio between mode number $M$ and photon number $N$. Generally, most of the current theoretical Boson Sampling hardness arguments (including our current work) consider only collision-free outcomes and rule out collision outcomes due to the convenience of mathematical analysis [aaronson2011computational, deshpande2022quantum, grier2022complexity, bouland2022noise, bouland2024average]. This calls for the “dilute" regime $M=\Omega(N^{2})$ to ensure the dominance of collision-free outcomes over the probability weight. In contrast, all the experiments to date [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian] operate in the “saturated regime” where the number of modes is comparable to the number of photons. Therefore, to reduce this gap, extending our hardness result of shallow-depth Boson Sampling to the saturated regime $M=O(N^{\gamma})$ for $1\leq\gamma<2$ is crucial, similarly to [bouland2023complexity] for the general Boson Sampling case. To do so, one also needs to consider collision outcomes in the analysis because collision outcomes now dominate the probability weight in this regime. Here, given that our worst-case hardness result in Theorem 3 holds for a linear number of modes $M=\Omega(N)$, this opens the possibility of establishing average-case hardness in the saturated regime by extending the worst-to-average-case reduction from Theorem 3 to include collision outcomes. However, since the collision outcomes do not interchange with each other generally by permutations (in contrast to collision-free outcomes), our proof technique in Sec. 4 using random permutation does not work for general collision outcomes. Therefore, more advanced proof techniques are required to address this challenge. In Appendix H, we present an approach that could offer a useful insight toward extending our average-case hardness result to include collision outcomes. Also, another possible way is to analytically prove that the local random circuit for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ is close enough to those of global Haar random circuits (numerically shown in [go2024exploring] and Appendix I). If this can be done, we can employ the previous argument [bouland2023complexity], which used global Haar random circuits, for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ in the saturated regime.
@@ Appendix C Proof of Theorem 3 (worst-case hardness)
We prove Lemma 7 by showing that the constant-depth linear optical circuit $C_{0}$ in Lemma 5 can be embedded into a $3$-dimensional grid-structured circuit, which can be implemented in the $\mathcal{B}\mathcal{B}^{*}$ architecture under a suitable permutation of mode indices. For interested readers, we provide the full proof of Lemma 7 in Appendix D.
@@ D.2 Encoding a grid linear optical circuit to the butterfly circuit architecture $\mathcal{B}$
Suppose that a mode number is given by $M=2^{n}$ for $n\in\mathbb{Z}^{+}$. For an $M$-mode single parallel $1$-dimensional grid linear optical circuit $C$, there exist an $M$-mode circuit $C^{\prime}$ in $\mathcal{B}$ and an $M$-mode permutation matrix $\bm{P}$ that satisfy $C=\bm{P}C^{\prime}\bm{P}^{T}$